# Revisiting Graph Analytics Benchmark

### v1

@@ Abstract.
The rise of graph analytics platforms has led to the development of various benchmarks for evaluating and comparing platform performance. However, existing benchmarks often fall short of fully assessing performance due to limitations in core algorithm selection, data generation processes (and the corresponding synthetic datasets), as well as the neglect of API usability evaluation. To address these shortcomings, we propose a novel graph analytics benchmark. First, we select eight core algorithms by extensively reviewing both academic and industrial settings. Second, we design an efficient and flexible data generator and produce eight new synthetic datasets as the default datasets for our benchmark. Lastly, we introduce a multi-level large language model (LLM)-based framework for API usability evaluation—the first of its kind in graph analytics benchmarks. We conduct comprehensive experimental evaluations on existing platforms ($\mathsf{GraphX}$, $\mathsf{PowerGraph}$, $\mathsf{Flash}$, $\mathsf{Grape}$, $\mathsf{Pregel+}$, $\mathsf{Ligra}$, and $\mathsf{G}$-$\mathsf{thinker}$). The experimental results demonstrate the superiority of our proposed benchmark.
Graph Analytics Benchmarks, Data Generator, API Usability Evaluation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Database performance evaluation
@@ 1. Introduction
Graph analytics platforms, such as $\mathsf{GraphX}$ (Gonzalez et al., 2014), $\mathsf{PowerGraph}$ (Gonzalez et al., 2012), $\mathsf{Flash}$ (Li et al., 2023a), $\mathsf{Grape}$ (Fan et al., 2018), $\mathsf{Pregel+}$ (Yan et al., 2015, 2014b; Lu et al., 2014), $\mathsf{Ligra}$ (Shun and Blelloch, 2013a), and $\mathsf{G}$-$\mathsf{thinker}$ (Yan et al., 2020), have been developed to facilitate large-scale graph data analysis. These platforms vary significantly in their design principles, architectures, and implementation details, which raises an important and practical question: how to evaluate graph analytics platforms and provide guidance on selecting the most appropriate one for specific environments.
@@
Accordingly, graph analytics benchmarks like $\mathsf{LDBC\leavevmode\nobreak\ Graphalytics}$ (Iosup et al., 2016) and $\mathsf{Graph500}$ (Graph500 Executive Committee, 2010) are commonly used to evaluate platform performance. These benchmarks provide a standardized framework for assessment, typically consisting of two main components: core algorithms and datasets. Core algorithms measure the efficiency of various graph processing tasks, offering a consistent basis for comparing platforms. Datasets aim to represent diverse graph characteristics found in real-world applications. However, due to high collection costs and privacy concerns, real-world datasets are often limited. As a result, benchmarks frequently use data generators to produce synthetic datasets that mimic real-world graphs.
@@
Motivation. Considering the goals and key components of graph analytics benchmarks, an ideal benchmark should select the minimal number of algorithms while ensuring they are representative and diverse enough to effectively distinguish the performance of various graph analytics platforms. The data generator should also efficiently produce synthetic datasets, allowing users to control input parameters to generate graphs with diverse characteristics. Unfortunately, existing benchmarks fall short of these ideals. Take the most popular $\mathsf{LDBC\leavevmode\nobreak\ Graphalytics}$ benchmark (Iosup et al., 2016) as an example. Regarding core algorithms, $\mathsf{LDBC\leavevmode\nobreak\ Graphalytics}$ selects PageRank ($\mathsf{PR}$(Page et al., 1999), Breadth First Search ($\mathsf{BFS}$(Lee, 1961), Single Source Shortest Path ($\mathsf{SSSP}$(Dijkstra, 1959), Weakly Connected Component ($\mathsf{WCC}$(Tarjan and Van Leeuwen, 1984), Label Propagation Algorithm ($\mathsf{LPA}$(Raghavan et al., 2007), Local Clustering Coefficient ($\mathsf{LCC}$(Holland and Leinhardt, 1971) as its core algorithms. However, as analyzed in Section 3, these six algorithms lack diversity, with most having linear time complexity, limiting their ability to reveal platform performance bottlenecks. Regarding dataset generation, $\mathsf{LDBC\leavevmode\nobreak\ Graphalytics}$’s data generator employs a sampling-based strategy to generate datasets, which, as examined in Section 4, suffers from efficiency issues due to numerous failed trials. Additionally, the generator only allows control over the scale of datasets, lacking flexibility in mimicking real-world datasets of different characteristics.
@@
Moreover, current benchmarks predominantly assess and compare performance metrics such as execution time, throughput, and scalability. However, these benchmarks often overlook the usability of the application programming interfaces (APIs) provided by graph analytics platforms. APIs play a crucial role in enhancing developer productivity, minimizing errors, and fostering wider platform adoption. The success of a graph analytics platform is significantly tied to the usability of its APIs, as highlighted in studies (Rauf et al., 2019; Myers, 2017). Consequently, API usability should be a key component in the design of graph analytics benchmarks. Yet, existing benchmarks typically neglect API usability, relying instead on simplistic measures like the number of lines of code to implement tasks (Zhu et al., 2014; Nam et al., 2019; Watson, 2009). Such metrics are insufficient for a comprehensive evaluation of API usability.
@@
Our Approach. We address the three major limitations in existing graph analytics benchmarks. For the core algorithms, we conduct an extensive review of both academic research and industrial applications, selecting eight algorithms ($\mathsf{PR}$, $\mathsf{SSSP}$, $\mathsf{LPA}$, $\mathsf{WCC}$, Triangle Counting ($\mathsf{TC}$), Betweenness Centrality ($\mathsf{BC}$), $k$-Clique ($\mathsf{KC}$), Core Decomposition ($\mathsf{CD}$)) based on their popularity, time complexity, and diversity of algorithmic topics. For the data generator, we redesign the edge sampling strategy and introduce a failure-free generator that eliminates the inefficiencies caused by failed trials in $\mathsf{LDBC\leavevmode\nobreak\ Graphalytics}$. Additionally, recognizing the critical impact of graph density and diameter on platform performance (Yan et al., 2015; Fan et al., 2018; Li et al., 2023a), we develop a mechanism that allows users to control these two attributes of the generated datasets, significantly enhancing the flexibility of the data generator. For API usability evaluation, we harness the capabilities of large language models ($\mathsf{LLM}$s) and propose a multi-level $\mathsf{LLM}$-based evaluation framework that simulates programmers of varying expertise levels, providing a comprehensive assessment of API usability. Table 1 compares our proposed benchmark with existing graph analytics benchmarks regarding core algorithms, controllable attributes of the synthetic datasets, and the evaluation metrics.
@@
$\mathsf{Graph500}$ (Graph500 Executive Committee, 2010) $\mathsf{BFS}$, $\mathsf{SSSP}$ Scale Timing, Throughput
$\mathsf{WGB}$ (Ammar and Özsu, 2013) $\mathsf{K}$-$\mathsf{Hop}$ (Wang et al., 2009), $\mathsf{SSSP}$, $\mathsf{PR}$, $\mathsf{WCC}$, $\mathsf{Cluster}$ (Schaeffer, 2007) Scale, Density Timing, Scalability
$\mathsf{BigDataBench}$ (Wang et al., 2014) $\mathsf{BFS}$, $\mathsf{PR}$, $\mathsf{WCC}$, $\mathsf{Cluster}$ Scale Timing, Throughput, MIPS, cache MPKI
$\mathsf{LDBC\leavevmode\nobreak\ Graphalytics}$ (Iosup et al., 2016) $\mathsf{PR}$, $\mathsf{BFS}$, $\mathsf{SSSP}$, $\mathsf{WCC}$, $\mathsf{LPA}$, $\mathsf{LCC}$ Scale Timing, Throughput, Scalability, Robustness
@@
  • We select eight representative core algorithms, spanning a wide range of algorithm popularity, algorithm diversity and computing models, to ensure comprehensive and realistic evaluation of platform performance (Section 3).
@@
  • We propose a $\mathsf{LLM}$-based evaluation framework to assess the API usability. To the best of our knowledge, this is the first graph analytics benchmark that includes a rigorous method for evaluating API usability (Section 5).
@@ 2. Related Work
Benchmarks for Graph Analytics Platforms. Benchmarking graph analytics platforms has attracted lots of attention in recent years, and many general-purpose benchmarks have been proposed in the literature. $\mathsf{Graph500}$ (Graph500 Executive Committee, 2010) uses $\mathsf{BFS}$ and $\mathsf{SSSP}$ algorithms on synthetic datasets generated by the $\mathsf{Kronecker}$ graph generator to measure platform performance in terms of timing and throughput. $\mathsf{WGB}$ (Ammar and Özsu, 2013) provides an efficient data generator for creating dynamic graphs similar to real-world graphs, offering a universal benchmark for graph analytics platforms. $\mathsf{BigDataBench}$ (Wang et al., 2014) aims to provide a comprehensive benchmarking suite for data systems, encompassing various data types, including graph data, text, and tables, and focusing on performance, energy efficiency, and cost-effectiveness under diverse workloads and data scenarios. $\mathsf{LDBC}$ (Iosup et al., 2016) is the most popular benchmark for graph analysis systems, offering datasets and evaluation metrics for various tasks, establishing itself as the state-of-the-art tool for assessing graph analytics platforms.
@@
Synthetic Graph Data Generators in Benchmarks. Graph data generator creates synthetic graph data that simulates real-world network structures and behaviors for use in benchmarking and analysis. Traditional Erdős-Rényi generator (Erdos et al., 1960) iteratively selects two random vertices to generate an edge. However, the generated graph has a low clustering coefficient and follows a Poisson degree distribution, which is contrary to real-world networks. Two improved generators, $\mathsf{Watts}$-$\mathsf{Strogatz}$ (Watts and Strogatz, 1998) and Barabási-$\mathsf{Albert}$ (Barabási and Albert, 1999), can generate highly-clustered and power-law graphs, respectively, while the $\mathsf{Kronecker}$ (Leskovec et al., 2010) generator used in $\mathsf{Graph500}$  (Murphy et al., 2010) can simultaneously follow these two properties. The $\mathsf{LFR3}$ (Lancichinetti et al., 2008) generator allows for the tuning of parameters to control clusters and power-law distribution. The $\mathsf{LDBC}$-$\mathsf{DG}$ generator can further generate and fit arbitrary degree distribution and has a better community similarity (Erling et al., 2015). In addition to the general graph generator, the $\mathsf{WGB}$ benchmark (Ammar and Özsu, 2013) adds a dynamic graph generator for evaluating systems under dynamic conditions. Many generators are integrated in graph libraries such as $\mathsf{NetworkX}$ (Hagberg et al., 2008) and $\mathsf{Neo4j}$-$\mathsf{APOC}$ (neo, [n. d.]). However, these generators are not efficient enough and cannot control the density and diameter of the generated graphs.
@@
API Usability Evaluation. Due to the importance of APIs in software development, many studies (Rauf et al., 2019; Myers, 2017; Piccioni et al., 2013a; Grill et al., 2012; Mosqueira-Rey et al., 2018; Farooq et al., 2010) have explored methods for evaluating API usability to guide API design. Existing works mainly rely on expert evaluations by project managers, developers, and testers, who review the API design and documentation to evaluate its usability and suggest improvements (Rauf et al., 2019; Farooq et al., 2010), while others gather insights through long-term tracking of developer feedback (Mosqueira-Rey et al., 2018). Brad A. Myers (Myers, 2017) points out that better API usability requires designing APIs to effectively meet the needs of users at different levels, including novices, professionals, and end-user programmers. Therefore, some studies (Myers, 2017; Piccioni et al., 2013a) have adopted the method of collecting feedback from users with varying levels of expertise to evaluate API usability. Clearly, this approach is costly and difficult to scale for large benchmark testing.
@@ 3. Core Algorithms
Algorithms play a vital role in benchmarking distributed graph processing platforms. $\mathsf{LDBC}$ (Iosup et al., 2016), the most popular benchmark for graph processing platforms, selects six algorithms ($\mathsf{BFS}$, $\mathsf{PR}$, $\mathsf{WCC}$, $\mathsf{LPA}$, $\mathsf{LCC}$ and $\mathsf{SSSP}$) frequently mentioned in academic papers as its core algorithms. However, the $\mathsf{LDBC}$’s algorithm set lacks diversity, as most of the algorithms focus on community detection and traversal, ignoring the suitability to different computing models.
To overcome these limitations, we consider multiple criteria when selecting core algorithms to ensure a comprehensive and representative benchmark including (1) popularity in academic research and real-world application, (2) diversity in algorithm topics, and (3) suitability for distributed graph computing models. Based on these criteria, we select eight algorithms, including:

### v2

@@ Abstract.
The rise of graph analytics platforms has led to the development of various benchmarks for evaluating and comparing platform performance. However, existing benchmarks often fall short of fully assessing performance due to limitations in core algorithm selection, data generation processes (and the corresponding synthetic datasets), as well as the neglect of API usability evaluation. To address these shortcomings, we propose a novel graph analytics benchmark. First, we select eight core algorithms by extensively reviewing both academic and industrial settings. Second, we design an efficient and flexible data generator and produce eight new synthetic datasets as the default datasets for our benchmark. Lastly, we introduce a multi-level large language model (LLM)-based framework for API usability evaluation-the first of its kind in graph analytics benchmarks. We conduct comprehensive experimental evaluations on existing platforms ($\mathsf{GraphX}$, $\mathsf{PowerGraph}$, $\mathsf{Flash}$, $\mathsf{Grape}$, $\mathsf{Pregel+}$, $\mathsf{Ligra}$, and $\mathsf{G}$-$\mathsf{thinker}$). The experimental results demonstrate the superiority of our proposed benchmark.
Graph Analytics Benchmarks; Data Generator; API Usability Evaluation
copyright: acmlicensedjournal: PACMMODjournalyear: 2025journalvolume: 3journalnumber: 3 (SIGMOD)article: 208publicationmonth: 6doi: 10.1145/3725345ccs: Information systems Database performance evaluation
@@ 1. Introduction
Graph analytics platforms, such as $\mathsf{GraphX}$ (gonzalez2014graphx), $\mathsf{PowerGraph}$ (gonzalez2012powergraph), $\mathsf{Flash}$ (li2023flash), $\mathsf{Grape}$ (fan2018parallelizing), $\mathsf{Pregel+}$ (yan2015effective; yan2014pregel; lu2014large), $\mathsf{Ligra}$ (shun2013ligra), and $\mathsf{G}$-$\mathsf{thinker}$ (yan2020g), have been developed to facilitate large-scale graph data analysis. These platforms vary significantly in their design principles, architectures, and implementation details, which raises an important and practical question: how to evaluate graph analytics platforms and provide guidance on selecting the most appropriate one for specific environments.
@@
Accordingly, graph analytics benchmarks like $\mathsf{LDBC\penalty 10000\ Graphalytics}$ (DBLP:journals/pvldb/IosupHNHPMCCSAT16) and $\mathsf{Graph500}$ (graph500) are commonly used to evaluate platform performance. These benchmarks provide a standardized framework for assessment, typically consisting of two main components: core algorithms and datasets. Core algorithms measure the efficiency of various graph processing tasks, offering a consistent basis for comparing platforms. Datasets aim to represent diverse graph characteristics found in real-world applications. However, due to high collection costs and privacy concerns, real-world datasets are often limited. As a result, benchmarks frequently use data generators to produce synthetic datasets that mimic real-world graphs.
@@
Motivation. Considering the goals and key components of graph analytics benchmarks, an ideal benchmark should select the minimal number of algorithms while ensuring they are representative and diverse enough to effectively distinguish the performance of various graph analytics platforms. The data generator should also efficiently produce synthetic datasets, allowing users to control input parameters to generate graphs with diverse characteristics. Unfortunately, existing benchmarks fall short of these ideals. Take the most popular $\mathsf{LDBC\penalty 10000\ Graphalytics}$ benchmark (DBLP:journals/pvldb/IosupHNHPMCCSAT16) as an example. Regarding core algorithms, $\mathsf{LDBC\penalty 10000\ Graphalytics}$ selects PageRank ($\mathsf{PR}$) (page1999pagerank), Breadth First Search ($\mathsf{BFS}$) (lee1961algorithm), Single Source Shortest Path ($\mathsf{SSSP}$) (DBLP:journals/nm/Dijkstra59), Weakly Connected Component ($\mathsf{WCC}$) (tarjan1984worst), Label Propagation Algorithm ($\mathsf{LPA}$) (raghavan2007near), Local Clustering Coefficient ($\mathsf{LCC}$) (holland1971transitivity) as its core algorithms. However, as analyzed in Section 3, these six algorithms lack diversity, with most having linear time complexity, limiting their ability to reveal platform performance bottlenecks. Regarding dataset generation, $\mathsf{LDBC\penalty 10000\ Graphalytics}$’s data generator employs a sampling-based strategy to generate datasets, which, as examined in Section 4, suffers from efficiency issues due to numerous failed trials. Additionally, the generator only allows control over the scale of datasets, lacking flexibility in mimicking real-world datasets of different characteristics.
@@
Moreover, current benchmarks predominantly assess and compare performance metrics such as execution time, throughput, and scalability. However, these benchmarks often overlook the usability of the application programming interfaces (APIs) provided by graph analytics platforms. APIs play a crucial role in enhancing developer productivity, minimizing errors, and fostering wider platform adoption. The success of a graph analytics platform is significantly tied to the usability of its APIs, as highlighted in studies (DBLP:journals/csr/RaufTP19; DBLP:conf/icse/Myers17). Consequently, API usability should be a key component in the design of graph analytics benchmarks. Yet, existing benchmarks typically neglect API usability, relying instead on simplistic measures like the number of lines of code to implement tasks (6976096; DBLP:conf/kbse/NamHMMV19; watson2009improving). Such metrics are insufficient for a comprehensive evaluation of API usability.
@@
Our Approach. We address the three major limitations in existing graph analytics benchmarks. For the core algorithms, we conduct an extensive review of both academic research and industrial applications, selecting eight algorithms ($\mathsf{PR}$, $\mathsf{SSSP}$, $\mathsf{LPA}$, $\mathsf{WCC}$, Triangle Counting ($\mathsf{TC}$), Betweenness Centrality ($\mathsf{BC}$), $k$-Clique ($\mathsf{KC}$), Core Decomposition ($\mathsf{CD}$)) based on their popularity, time complexity, and diversity of algorithmic topics. For the data generator, we redesign the edge sampling strategy and introduce a failure-free generator that eliminates the inefficiencies caused by failed trials in $\mathsf{LDBC\penalty 10000\ Graphalytics}$. Additionally, recognizing the critical impact of graph density and diameter on platform performance (yan2015effective; fan2018parallelizing; li2023flash), we develop a mechanism that allows users to control these two attributes of the generated datasets, significantly enhancing the flexibility of the data generator. For API usability evaluation, we harness the capabilities of large language models ($\mathsf{LLM}$s) and propose a multi-level $\mathsf{LLM}$-based evaluation framework that simulates programmers of varying expertise levels, providing a comprehensive assessment of API usability. Table 1 compares our proposed benchmark with existing graph analytics benchmarks regarding core algorithms, controllable attributes of the synthetic datasets, and the evaluation metrics.
@@
$\mathsf{Graph500}$ (graph500) $\mathsf{BFS}$, $\mathsf{SSSP}$ Scale Timing, Throughput
$\mathsf{WGB}$ (DBLP:conf/wbdb/AmmarO13) $\mathsf{K}$-$\mathsf{Hop}$ (wang2009mobility), $\mathsf{SSSP}$, $\mathsf{PR}$, $\mathsf{WCC}$, $\mathsf{Cluster}$ (schaeffer2007graph) Scale, Density Timing, Scalability
$\mathsf{BigDataBench}$ (DBLP:conf/hpca/WangZLZYHGJSZZLZLQ14) $\mathsf{BFS}$, $\mathsf{PR}$, $\mathsf{WCC}$, $\mathsf{Cluster}$ Scale Timing, Throughput, MIPS, cache MPKI
$\mathsf{LDBC\penalty 10000\ Graphalytics}$ (DBLP:journals/pvldb/IosupHNHPMCCSAT16) $\mathsf{PR}$, $\mathsf{BFS}$, $\mathsf{SSSP}$, $\mathsf{WCC}$, $\mathsf{LPA}$, $\mathsf{LCC}$ Scale Timing, Throughput, Scalability, Robustness
@@
  • We select eight representative core algorithms, spanning a wide range of algorithm popularity, algorithm diversity and computing models, to ensure a comprehensive and realistic evaluation of platform performance (Section 3).
@@
  • We propose an $\mathsf{LLM}$-based evaluation framework to assess the API usability. To the best of our knowledge, this is the first graph analytics benchmark that includes a rigorous method for evaluating API usability (Section 5).
@@ 2. Related Work
Benchmarks for Graph Analytics Platforms. Benchmarking graph analytics platforms has attracted lots of attention in recent years, and many general-purpose benchmarks have been proposed in the literature. $\mathsf{Graph500}$ (graph500) uses $\mathsf{BFS}$ and $\mathsf{SSSP}$ algorithms on synthetic datasets generated by the $\mathsf{Kronecker}$ graph generator to measure platform performance in terms of timing and throughput. $\mathsf{WGB}$ (DBLP:conf/wbdb/AmmarO13) provides an efficient data generator for creating dynamic graphs similar to real-world graphs, offering a universal benchmark for graph analytics platforms. $\mathsf{BigDataBench}$ (DBLP:conf/hpca/WangZLZYHGJSZZLZLQ14) aims to provide a comprehensive benchmarking suite for data systems, encompassing various data types, including graph data, text, and tables, and focusing on performance, energy efficiency, and cost-effectiveness under diverse workloads and data scenarios. $\mathsf{LDBC}$ (DBLP:journals/pvldb/IosupHNHPMCCSAT16) is the most popular benchmark for graph analysis systems, offering datasets and evaluation metrics for various tasks, establishing itself as the state-of-the-art tool for assessing graph analytics platforms.
@@
Synthetic Graph Data Generators in Benchmarks. The graph data generator creates synthetic graph data that simulates real-world network structures and behaviors for use in benchmarking and analysis. Traditional Erdős-Rényi generator (erdos1960evolution) iteratively selects two random vertices to generate an edge. However, the generated graph has a low clustering coefficient and follows a Poisson degree distribution, which is contrary to real-world networks. Two improved generators, $\mathsf{Watts}$-$\mathsf{Strogatz}$ (watts1998collective) and Barabási-$\mathsf{Albert}$ (barabasi1999emergence), can generate highly-clustered and power-law graphs, respectively, while the $\mathsf{Kronecker}$ (leskovec2010kronecker) generator used in $\mathsf{Graph500}$ (murphy2010introducing) can simultaneously follow these two properties. The $\mathsf{LFR3}$ (lancichinetti2008benchmark) generator allows for the tuning of parameters to control clusters and power-law distribution. The $\mathsf{LDBC}$-$\mathsf{DG}$ generator can generate graphs with arbitrary degree distributions and better community similarity (erling2015ldbc). In addition to the general graph generator, the $\mathsf{WGB}$ benchmark (DBLP:conf/wbdb/AmmarO13) adds a dynamic graph generator for evaluating systems under dynamic conditions. Many generators are integrated in graph libraries such as $\mathsf{NetworkX}$ (hagberg2008exploring) and $\mathsf{Neo4j}$-$\mathsf{APOC}$ (neo4j-apoc). However, these generators are not efficient enough and cannot control the density and diameter of the generated graphs.
@@
API Usability Evaluation. Due to the importance of APIs in software development, many studies (DBLP:journals/csr/RaufTP19; DBLP:conf/icse/Myers17; DBLP:conf/esem/PiccioniFM13; DBLP:conf/hcse/GrillPT12; DBLP:journals/infsof/Mosqueira-ReyAM18; DBLP:conf/chi/FarooqWZ10) have explored methods for evaluating API usability to guide API design. Existing works mainly rely on expert evaluations by project managers, developers, and testers, who review the API design and documentation to evaluate its usability and suggest improvements (DBLP:journals/csr/RaufTP19; DBLP:conf/chi/FarooqWZ10), while others gather insights through long-term tracking of developer feedback (mosqueira2018systematic). Brad A. Myers (DBLP:conf/icse/Myers17) points out that better API usability requires designing APIs to effectively meet the needs of users at different levels, including novices, professionals, and end-user programmers. Therefore, some studies (DBLP:conf/icse/Myers17; DBLP:conf/esem/PiccioniFM13) have adopted the method of collecting feedback from users with varying levels of expertise to evaluate API usability. Clearly, this approach is costly and difficult to scale for large benchmark testing.
@@ 3. Core Algorithms
Algorithms play a vital role in benchmarking distributed graph processing platforms. $\mathsf{LDBC}$ (DBLP:journals/pvldb/IosupHNHPMCCSAT16), the most popular benchmark for graph processing platforms, selects six algorithms ($\mathsf{BFS}$, $\mathsf{PR}$, $\mathsf{WCC}$, $\mathsf{LPA}$, $\mathsf{LCC}$ and $\mathsf{SSSP}$) frequently mentioned in academic papers as its core algorithms. However, $\mathsf{LDBC}$’s algorithm set lacks diversity, as most of the algorithms focus on community detection and traversal, while ignoring the suitability for different computing models.
To overcome these limitations, we consider multiple criteria when selecting core algorithms to ensure a comprehensive and representative benchmark, including (1) popularity in academic research and real-world application, (2) diversity in algorithm topics, and (3) suitability for distributed graph computing models. Based on these criteria, we select eight algorithms, including: